# Score Forgetting Distillation: A Swift, Data-Free Method for Machine Unlearning in Diffusion Models

### v2

@@ Abstract
The machine learning community is increasingly recognizing the importance of fostering trust and safety in modern generative AI (GenAI) models. We posit machine unlearning (MU) as a crucial foundation for developing safe, secure, and trustworthy GenAI models. Traditional MU methods often rely on stringent assumptions and require access to real data. This paper introduces Score Forgetting Distillation (SFD), an innovative MU approach that promotes the forgetting of undesirable information in diffusion models by aligning the conditional scores of “unsafe” classes or concepts with those of “safe” ones. To eliminate the need for real data, our SFD framework incorporates a score-based MU loss into the score distillation objective of a pretrained diffusion model. This serves as a regularization term that preserves desired generation capabilities while enabling the production of synthetic data through a one-step generator. Our experiments on pretrained label-conditional and text-to-image diffusion models demonstrate that our method effectively accelerates the forgetting of target classes or concepts during generation, while preserving the quality of other classes or concepts. This unlearned and distilled diffusion not only pioneers a novel concept in MU but also accelerates the generation speed of diffusion models. Our experiments and studies on a range of diffusion models and datasets confirm that our approach is generalizable, effective, and advantageous for MU in diffusion models.
@@ 1 Introduction
Diffusion models, also known as score-based generative models (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Dhariwal and Nichol, 2021; Karras et al., 2022), have emerged as the leading choice for generative modeling of high-dimensional data. These models are widely celebrated for their ability to produce high-quality, diverse, and photorealistic images (Nichol et al., 2022; Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Podell et al., 2024; Zheng et al., 2024). However, their capacity to memorize and reproduce specific images and concepts from training datasets raises significant privacy and safety concerns. Moreover, they are susceptible to poisoning attacks, enabling the generation of targeted images with embedded triggers, posing substantial security risks (Rando et al., 2022; Chen et al., 2023b).
@@
To address these challenges, we introduce Score Forgetting Distillation (SFD), a novel framework designed to efficiently mitigate the influence of specific characteristics in data points on pre-trained diffusion models. This framework is a key part of the broader domain of Machine Unlearning (MU), which has evolved significantly to address core issues in trustworthy machine learning (Lowd and Meek, 2005; Narayanan and Shmatikov, 2008; Abadi et al., 2016). Originating from compliance needs with data protection regulations such as the “right to be forgotten” (Hoofnagle et al., 2019), MU has broadened its scope to include applications in diffusion modeling across various domains like computer vision and content generation (Gandikota et al., 2023; Fan et al., 2024; Heng and Soh, 2024). Additionally, MU aims to promote model fairness (Oesterling et al., 2024), refine pre-training methodologies (Jain et al., 2023; Jia et al., 2023), and reduce the generation of inappropriate content (Gandikota et al., 2023). The development of SFD is aligned with these objectives, providing a strategic approach to mitigate the potential risks and reduce the high generation costs associated with diffusion models, thereby advancing the field of trustworthy machine learning.
@@
MU methods are generally categorized into two types: exact MU and approximate MU. Exact MU entails creating a model that behaves as if sensitive data had never been part of the training set (Cao and Yang, 2015; Bourtoule et al., 2021). This process requires the unlearned model to be identical in distribution to a model retrained without the sensitive data, both in terms of model weights and output behavior. In contrast, approximate MU does not seek an exact match between the unlearned model and a retrained model. Instead, it aims to approximate how closely the output distributions of the two models align after the unlearning process. A prominent strategy in approximate MU utilizes the principles of differential privacy (Dwork, 2006). For instance, Guo et al. (2019) introduced a certified removal technique that prevents adversaries from extracting information about removed training data, offering a theoretical guarantee of data privacy. However, these approaches typically necessitate retraining the model from scratch, which can be computationally intensive and require access to the original training dataset. Efficient and stable unlearning has become crucial in MU. Techniques like the influence functions (Warnecke et al., 2021; Izzo et al., 2021), selective forgetting (Golatkar et al., 2020), weight-based pruning (Liu et al., 2024), and gradient-based saliency (Fan et al., 2024) have been explored, though they often suffer from performance degradation or restrictive assumptions (Becker and Liebig, 2022). These methods are primarily applied to MU for image classification tasks and do not adequately address the rapid forgetting and unlearning required for data generation tasks.
@@
Given the prominence of diffusion models, there is a growing need to develop MU techniques that specifically cater to these models, ensuring efficient unlearning while maintaining generation capabilities (Gandikota et al., 2023; Fan et al., 2024; Heng and Soh, 2024). Our SFD framework efficiently distills the knowledge from a pre-trained diffusion model by optimizing two learnable modules—a generator network and a score network—guided by the frozen pre-trained model itself. The score network is trained to optimize the score associated with the generator by minimizing a score distillation loss, which aims to match the conditional scores of the class to forget and the classes to remember with those of the pre-trained model. The generator network learns to produce examples that are “indistinguishable” by the pre-trained score network and fake score network in terms of score predictions, utilizing a model-based cross-class score distillation loss.
@@ 2 Method
Diffusion models are celebrated for their superior performance in generating high-quality and diverse samples. However, their robust capabilities also introduce challenges, particularly the risk of misuse in generating inappropriate content. This concern highlights the ethical implications and potential negative impacts of their application. Additionally, these models have a significant drawback: slow sampling speeds. This inefficiency becomes particularly problematic in downstream tasks that require finetuning on synthetic data generated by these models. When access to real data is not feasible, the task of preparing a sufficiently large synthetic dataset can already become computationally prohibitive (Yin et al., 2024). This issue is especially acute in the context of MU and image generation, where access to real data often raises privacy concerns, making reliance on synthetic data crucial. Consequently, the slow sampling rate of diffusion models presents a critical bottleneck, necessitating improvements to enable effective data-free MU operations.
In this section, we introduce SFD, a principled and data-free approach designed to address the MU problem while simultaneously achieving fast sampling for diffusion models. Building on recent advancements in data-free diffusion distillation for one-step generation (Luo et al., 2023; Zhou et al., 2024b), we conceptualize MU in diffusion models as a problem of MU-regularized score distillation.
@@ 2.1 Problem Definition and Notations
Before diving into the specific MU problem, we will first establish the essential concepts and notations in diffusion modeling: A diffusion model corrupts its data $x\sim p_{\text{data}}(x\mid c)$ during the forward diffusion process at time $t$ as $z_{t}=a_{t}x+\sigma_{t}\epsilon_{t}$, where $\epsilon_{t}\sim\mathcal{N}(0,1)$, $c$ represents the given condition such as a label or text, and $a_{t}$ and $\sigma_{t}$ are diffusion scheduling parameters. The goal of pretraining a diffusion model is to obtain an optimal score estimator $s_{\phi}(z_{t},c,t)$ such that $s_{\phi}(z_{t},c,t)=\nabla_{z_{t}}\ln p_{\text{data}}(z_{t}\,|\,c)$. Let $x_{\phi}(z_{t},c,t)$ be the optimal conditional mean estimator such that for $x_{\phi}(z_{t},c,t)=\mathbb{E}[x\,|\,z_{t},c,t]$. Applying Tweedie’s formula (Robbins, 1992; Efron, 2011) in the context of diffusion modeling (Luo, 2022; Chung et al., 2023; Zhou et al., 2024b), the optimal score and conditional mean estimators, $s_{\phi}$ and $x_{\phi}$, for the training data are related as follows:
@@
With this optimal score estimator, we can construct a corresponding reverse diffusion process, enabling us to approximately sample from the data distribution through numerical discretization along the time horizon (Anderson, 1982; Song et al., 2020).
@@ 2.2 Score Forgetting Distillation
where $\omega_{t}>0$ is a re-weighting function, and $\psi^{*}(\theta)$ represents the optimal solution to the model-based explicit SM (MESM) loss, which can be expressed as
@@
where $\gamma_{t}>0$ is a re-weighting function. In practice, the lack of the access to $\nabla_{x}\ln p_{\theta}(z_{t}\,|\,c)$ makes Eq. 4 intractable. However, we can alternatively optimize a denoising SM loss (Vincent, 2011) as
@@
This formulation corresponds to a bi-level optimization problem (Ye et al., 1997; Hong et al., 2023; Shen et al., 2023), subject to an additional forgetting-based constraint. Solving this problem directly is challenging, so we initially relax the constraint specified by $\mathcal{L}_{\text{sfd}}$ in the above equation by integrating it into the distillation objective as an additional MU regularization term:
@@
where $\lambda$ and $\mu$ are tunable constants that serve as control knobs to balance the distillation of the remaining classes and the unlearning of the target class. Furthermore, we implement an alternating update strategy between $\theta$ and $\psi$. This approach mitigates the need to obtain the optimal score estimator $\psi^{*}(\theta)$ for each $\theta$, simplifying the computational process. We outline a practical implementation of this strategy in Algorithm 1. Specifically, generalizing the derivation in Zhou et al. (2024b), we have the following Lemma, whose proof is provided in Appendix D:
@@
A biased loss for $\theta$ can be derived by replacing $\psi^{*}(\theta)$ in either Eq. 3 or Eq. 8 with its SGD-based approximation $\psi$, and disregarding the dependency of $\psi^{*}$ on $\theta$ when computing the gradient of $\theta$. Empirical experiments by Zhou et al. (2024b) suggest that in the context of diffusion distillation without involving unlearning, Eq. 8 can be effective independently, while Eq. 3 may not perform as expected. This observation leads to a practical approach that involves subtracting Eq. 3 from Eq. 8. This strategy aims to sidestep detrimental biased gradient directions and potentially compensate for the overlooked gradient dependency of $\psi^{*}(\theta)$. We implement this approach in practice under the framework of SFD, defining the loss used in practice as follows:
@@
where $\alpha\geq 0$ is some constant that typically set as 1 or 1.2, $z_{t}=a_{t}x+\sigma_{t}{\epsilon}_{t},~{}x\sim\mathcal{D}_{\theta,c_{2}},~{}{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),~{}t\sim\text{Unif}[t_{\text{min}},t_{\text{max}}]$. In this paper, we follow Yin et al. (2024) and Zhou et al. (2024b) to set $\omega_{t}=\frac{\sigma_{t}^{4}}{a_{t}^{2}}\frac{C}{\|x_{\phi}(z_{t},t,c)-x\|_{1,\text{sg}}}$, where $C$ is the data dimension and “sg” stands for stop gradient.
@@ 3 Experiments
In our experiments, we thoroughly evaluate our method for class forgetting in diffusion models pretrained on two datasets, CIFAR-10 and STL-10, which have been commonly used for evaluating MU in previous studies. We provide the details of them in Appendix B. We also assess our method for concept forgetting tasks, such as celebrity forgetting, in text-to-image diffusion models.
@@
We explore class forgetting in class-conditional image generation tasks using DDPM (Ho et al., 2020), and investigate concept forgetting in text-to-image generation tasks using Stable Diffusion (Rombach et al., 2022). Class forgetting aims to prevent class-conditional diffusion models from generating images of a specified class, while concept forgetting seeks to remove the model’s ability to generate images containing specific concepts, such as celebrities or inappropriate content. Class-conditional and text-to-image sampling are achieved by inputting class labels and text prompts into the respective diffusion models, with fidelity further enhanced by classifier-free guidance introduced in Ho and Salimans (2022). Specifically, we approach unlearning by overriding a class or concept with another that is safe to retain. The class forgetting experiments were conducted on class-conditional diffusion models pre-trained on CIFAR-10 and STL-10, while the concept forgetting experiments were conducted on Stable Diffusion, including forgetting celebrities, specifically American actor Brad Pitt and actress Angelina Jolie, and forgetting a general NSFW (not safe for work) concept, $i.e.$, nudity. For DDPM baselines, we used the default 1000-step DDPM samplers to obtain FIDs for samples from the remaining classes, while for SD baselines, we used DDIM samplers with 50 steps. In contrast, our method requires only a single step for generation, making it 1,000 times faster than the DDPM baselines and 50 times faster in latent sampling than the SD baselines.
@@
To quantitatively assess the effectiveness of class forgetting, we primarily focus on the success rate of forgetting the target class, and the generative capability on classes to retain. Specifically, we measure the success rate of forgetting by unlearning accuracy (UA) employing an external classifier trained on the original training set, which is essentially the mis-classification rate of the classifier on the generated samples from the target class. We measure image generation quality using Fréchet Inception Distance (FID) (Heusel et al., 2017) and sample diversity using Inception Score (IS) (Salimans et al., 2016). Additionally, we report Precision and Recall (Kynkäänniemi et al., 2019), and number of function evaluations (NFEs) for sampling. Following Fan et al. (2024), we compute and report generation quality metrics using generated samples, with the full training set from the remaining classes serving as the reference. For concept forgetting tasks including celebrity forgetting and “nudity” forgetting, we also provide quantitative evaluations as well as qualitative comparison. Specifically, we evaluate celebrity forgetting using a off-the-shelf celebrity face detector, while we assess the MU performance of our “nudity” forgetting model on the I2P benchmark (https://github.com/ml-research/i2p). Please refer to Section B.2 for more details of the evaluation metrics.
@@
Our main implementation of class forgetting experiments is based on DDPM (Ho et al., 2020), where we utilize the codebase developed by Fan et al. (2024). Additionally, we implement our method using EDM (Karras et al., 2022) framework and the official codebase (https://github.com/NVlabs/edm). For concept forgetting experiments, we implement our method for SD models based on the implementation of Zhou et al. (2024a). We adopt the same model configuration for both the generator $g_{\theta}$ and its score estimation network $s_{\psi}$ and initialize the model weights according to the pre-trained score network $s_{\phi}$. This type of initialization prepares a good starting point for SFD.

### v3

@@ Abstract
The machine learning community is increasingly recognizing the importance of fostering trust and safety in modern generative AI (GenAI) models. We posit machine unlearning (MU) as a crucial foundation for developing safe, secure, and trustworthy GenAI models. Traditional MU methods often rely on stringent assumptions and require access to real data. This paper introduces Score Forgetting Distillation (SFD), an innovative MU approach that promotes the forgetting of undesirable information in diffusion models by aligning the conditional scores of “unsafe” classes or concepts with those of “safe” ones. To eliminate the need for real data, our SFD framework incorporates a score-based MU loss into the score distillation objective of a pretrained diffusion model. This serves as a regularization term that preserves desired generation capabilities while enabling the production of synthetic data through a one-step generator. Our experiments on pretrained label-conditional and text-to-image diffusion models demonstrate that our method effectively accelerates the forgetting of target classes or concepts during generation, while preserving the quality of other classes or concepts. This unlearned and distilled diffusion not only pioneers a novel concept in MU but also accelerates the generation speed of diffusion models. Our experiments and studies on a range of diffusion models and datasets confirm that our approach is generalizable, effective, and advantageous for MU in diffusion models. PyTorch code is available at https://github.com/tqch/score-forgetting-distillation.
@@ 1 Introduction
Diffusion models, also known as score-based generative models [69; 71; 31; 16; 39], have emerged as the leading choice for generative modeling of high-dimensional data. These models are widely celebrated for their ability to produce high-quality, diverse, and photorealistic images [54; 58; 62; 61; 56; 93]. However, their capacity to memorize and reproduce specific images and concepts from training datasets raises significant privacy and safety concerns. Moreover, they are susceptible to poisoning attacks, enabling the generation of targeted images with embedded triggers, posing substantial security risks [59; 11].
@@
To address these challenges, we introduce Score Forgetting Distillation (SFD), a novel framework designed to efficiently mitigate the influence of specific characteristics in data points on pre-trained diffusion models. This framework is a key part of the broader domain of Machine Unlearning (MU), which has evolved significantly to address core issues in trustworthy machine learning [45; 49; 1]. Originating from compliance needs with data protection regulations such as the “right to be forgotten” [33], MU has broadened its scope to include applications in diffusion modeling across various domains like computer vision and content generation [20; 19; 27]. Additionally, MU aims to promote model fairness [55], refine pre-training methodologies [36; 38], and reduce the generation of inappropriate content [20]. The development of SFD is aligned with these objectives, providing a strategic approach to mitigate the potential risks and reduce the high generation costs associated with diffusion models, thereby advancing the field of trustworthy machine learning.
@@
MU methods are generally categorized into two types: exact MU and approximate MU. Exact MU entails creating a model that behaves as if sensitive data had never been part of the training set [6; 5]. This process requires the unlearned model to be identical in distribution to a model retrained without the sensitive data, both in terms of model weights and output behavior. In contrast, approximate MU does not seek an exact match between the unlearned model and a retrained model. Instead, it aims to approximate how closely the output distributions of the two models align after the unlearning process. A prominent strategy in approximate MU utilizes the principles of differential privacy [17]. For instance, Guo et al. [24] introduced a certified removal technique that prevents adversaries from extracting information about removed training data, offering a theoretical guarantee of data privacy. However, these approaches typically necessitate retraining the model from scratch, which can be computationally intensive and require access to the original training dataset. Efficient and stable unlearning has become crucial in MU. Techniques like the influence functions [80; 35], selective forgetting [22], weight-based pruning [44], and gradient-based saliency [19] have been explored, though they often suffer from performance degradation or restrictive assumptions [4]. These methods are primarily applied to MU for image classification tasks and do not adequately address the rapid forgetting and unlearning required for data generation tasks.
@@
Given the prominence of diffusion models, there is a growing need to develop MU techniques that specifically cater to these models, ensuring efficient unlearning while maintaining generation capabilities [20; 19; 27]. Our SFD framework efficiently distills the knowledge from a pre-trained diffusion model by optimizing two learnable modules—a generator network and a score network—guided by the frozen pre-trained model itself. The score network is trained to optimize the score associated with the generator by minimizing a score distillation loss, which aims to match the conditional scores of the class to forget and the classes to remember with those of the pre-trained model. The generator network learns to produce examples that are “indistinguishable” by the pre-trained score network and fake score network in terms of score predictions, utilizing a model-based cross-class score distillation loss.
@@ 2 Method
Diffusion models are celebrated for their superior performance in generating high-quality and diverse samples. However, their robust capabilities also introduce challenges, particularly the risk of misuse in generating inappropriate content. This concern highlights the ethical implications and potential negative impacts of their application. Additionally, these models have a significant drawback: slow sampling speeds. This inefficiency becomes particularly problematic in downstream tasks that require finetuning on synthetic data generated by these models. When access to real data is not feasible, the task of preparing a sufficiently large synthetic dataset can already become computationally prohibitive [86]. This issue is especially acute in the context of MU and image generation, where access to real data often raises privacy concerns, making reliance on synthetic data crucial. Consequently, the slow sampling rate of diffusion models presents a critical bottleneck, necessitating improvements to enable effective data-free MU operations.
In this section, we introduce SFD, a principled and data-free approach designed to address the MU problem while simultaneously achieving fast sampling for diffusion models. Building on recent advancements in data-free diffusion distillation for one-step generation [47; 96], we conceptualize MU in diffusion models as a problem of MU-regularized score distillation.
@@ 2.1 Problem Definition and Notations
Before diving into the specific MU problem, we will first establish the essential concepts and notations in diffusion modeling: A diffusion model corrupts its data $x\sim p_{\text{data}}(x\mid c)$ during the forward diffusion process at time $t$ as $z_{t}=a_{t}x+\sigma_{t}\epsilon_{t}$, where $\epsilon_{t}\sim\mathcal{N}(0,1)$, $c$ represents the given condition such as a label or text, and $a_{t}$ and $\sigma_{t}$ are diffusion scheduling parameters. The goal of pretraining a diffusion model is to obtain an optimal score estimator $s_{\phi}(z_{t},c,t)$ such that $s_{\phi}(z_{t},c,t)=\nabla_{z_{t}}\ln p_{\text{data}}(z_{t}\,|\,c)$. Let $x_{\phi}(z_{t},c,t)$ be the optimal conditional mean estimator such that for $x_{\phi}(z_{t},c,t)=\mathbb{E}[x\,|\,z_{t},c,t]$. Applying Tweedie’s formula [60; 18] in the context of diffusion modeling [46; 13; 96], the optimal score and conditional mean estimators, $s_{\phi}$ and $x_{\phi}$, for the training data are related as follows:
@@
With this optimal score estimator, we can construct a corresponding reverse diffusion process, enabling us to approximately sample from the data distribution through numerical discretization along the time horizon [2; 72].
@@ 2.2 Score Forgetting Distillation
where $\omega_{t}>0$ is a re-weighting function, and $\psi^{*}(\theta)$ represents the optimal solution to the model-based explicit SM (MESM) loss, which is a Fisher divergence that can be expressed as
@@
where $\gamma_{t}>0$ is a re-weighting function. In practice, the lack of the access to $\nabla_{x}\ln p_{\theta}(z_{t}\,|\,c)$ makes Eq. 4 intractable. However, we can alternatively optimize a denoising SM loss [77] as
@@
This formulation corresponds to a bi-level optimization problem [84; 32; 68], subject to an additional forgetting-based constraint. Solving this problem directly is challenging, so we initially relax the constraint specified by $\mathcal{L}_{\text{sfd}}$ in the above equation by integrating it into the distillation objective as an additional MU regularization term:
@@
where $\lambda$ and $\mu$ are tunable constants that serve as control knobs to balance the distillation of the remaining classes and the unlearning of the target class. Furthermore, we implement an alternating update strategy between $\theta$ and $\psi$. This approach mitigates the need to obtain the optimal score estimator $\psi^{*}(\theta)$ for each $\theta$, simplifying the computational process. We outline a practical implementation of this strategy in Algorithm 1. Specifically, generalizing the derivation in Zhou et al. [96], we have the following Lemma, whose proof is provided in Appendix D:
@@
A biased loss for $\theta$ can be derived by replacing $\psi^{*}(\theta)$ in either Eq. 3 or Eq. 8 with its SGD-based approximation $\psi$, and disregarding the dependency of $\psi^{*}$ on $\theta$ when computing the gradient of $\theta$. Empirical experiments by Zhou et al. [96] suggest that in the context of diffusion distillation without involving unlearning, Eq. 8 can be effective independently, while Eq. 3 may not perform as expected. This observation leads to a practical approach that involves subtracting Eq. 3 from Eq. 8. This strategy aims to sidestep detrimental biased gradient directions and potentially compensate for the overlooked gradient dependency of $\psi^{*}(\theta)$. We implement this approach in practice under the framework of SFD, defining the loss used in practice as follows:
@@
where $\alpha\geq 0$ is some constant that is typically set as 1 or 1.2, $z_{t}=a_{t}x+\sigma_{t}{\epsilon}_{t},~{}x\sim\mathcal{D}_{\theta,c_{2}},~{}{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),~{}t\sim\text{Unif}[t_{\text{min}},t_{\text{max}}]$. In this paper, we follow Yin et al. [86] and Zhou et al. [96] to set $\omega_{t}=\frac{\sigma_{t}^{4}}{a_{t}^{2}}\frac{C}{\|x_{\phi}(z_{t},t,c)-x\|_{1,\text{sg}}}$, where $C$ is the data dimension and “sg” stands for stop gradient.
@@ 3 Experiments
In our experiments, we thoroughly evaluate our method for class forgetting in diffusion models pretrained on two datasets, CIFAR-10 and STL-10, which have been commonly used for evaluating MU in previous studies. We provide the details of them in Appendix B. We also assess our method for concept forgetting tasks, such as celebrity forgetting and “nudity” forgetting, in text-to-image diffusion models.
@@
We explore class forgetting in class-conditional image generation tasks using DDPM [31], and investigate concept forgetting in text-to-image generation tasks using Stable Diffusion (SD) [61]. Class forgetting aims to prevent class-conditional diffusion models from generating images of a specified class, while concept forgetting seeks to remove the model’s ability to generate images containing specific concepts, such as celebrities or inappropriate content. Class-conditional and text-to-image sampling are achieved by inputting class labels and text prompts into the respective diffusion models, with fidelity further enhanced by classifier-free guidance introduced in Ho & Salimans [30]. Specifically, we approach unlearning by overriding a class or concept with another that is safe to retain. The class forgetting experiments were conducted on class-conditional diffusion models pre-trained on CIFAR-10 and STL-10, while the concept forgetting experiments were conducted on Stable Diffusion, including forgetting celebrities, specifically American actor Brad Pitt and actress Angelina Jolie, and forgetting a general NSFW (not safe for work) concept, $i.e.$, nudity. For DDPM baselines, we used the default 1000-step DDPM samplers to obtain FIDs for samples from the remaining classes, while for SD baselines, we used DDIM samplers with 50 steps. In contrast, our method requires only a single step for generation, making it 1,000 times faster than the DDPM baselines and 50 times faster in latent sampling than the SD baselines.
@@
To quantitatively assess the effectiveness of class forgetting, we primarily focus on the success rate of forgetting the target class, and the generative capability on classes to retain. Specifically, we measure the success rate of forgetting by Unlearning Accuracy (UA) employing an external classifier trained on the original training set, which is essentially the mis-classification rate of the classifier on the generated samples from the target class. We measure image generation quality using Fréchet Inception distance (FID) [28] and sample diversity using Inception scores (IS) [64]. Additionally, we report Precision and Recall [43], and number of function evaluations (NFEs) for sampling. Following Fan et al. [19], we compute and report generation quality metrics using generated samples, with the full training set from the remaining classes serving as the reference. For concept forgetting tasks including celebrity forgetting and “nudity” forgetting, we also provide quantitative evaluations as well as qualitative comparison. Specifically, we evaluate celebrity forgetting using a off-the-shelf celebrity face detector, while we assess the MU performance of our “nudity” forgetting model on the I2P benchmark (https://github.com/ml-research/i2p). Please refer to Section B.2 for more details of the evaluation metrics.
@@
Our main implementation of class forgetting experiments is based on DDPM [31], where we utilize the codebase developed by Fan et al. [19]. Additionally, we implement our method using EDM [39] framework and the official codebase (https://github.com/NVlabs/edm). For concept forgetting experiments, we implement our method for SD models based on the implementation of Zhou et al. [95]. We adopt the same model configuration for both the generator $g_{\theta}$ and its score estimation network $s_{\psi}$ and initialize the model weights according to the pre-trained score network $s_{\phi}$. This type of initialization prepares a good starting point for SFD.